# DCoM: Active Learning for All Learners

### v2

@@ 3 Our method: DCoM
The error bound in (5) guides our method for active query selection, which aims to minimize this bound by maximizing the coverage $P(C({\mathbb{L}},\Delta))$ and individual purity $\pi(\delta_{i})=P(B_{\delta_{i}}(x_{i})\text{ is pure})$. Recall that ${\mathbb{L}}$ denotes the set of labeled points, $\delta_{i}$ the influence radius of $x_{i}\in{\mathbb{L}}$, $\Delta=[\delta_{i}]_{i=1}^{b}$, and
$C({\mathbb{L}},\Delta)=\bigcup_{(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta}B_{\delta_{i}}(x_{i}),\qquad\qquad\pi(\delta_{i})=P\left(\left\{x\in B_{\delta_{i}}\left(x_{i}\right):f(x)=f(x_{i})\right\}\right)$
Since it was shown in [39] that maximizing $P(C({\mathbb{L}},\delta))$ is already NP-hard, we opt for developing a greedy algorithm for actively selecting the query set, to be termed DCoM.
@@ 3.1 Active learning method
The algorithm involves several steps:
Initialization: Define a suitable embedding space for the data and initialize essential parameters.
If ${\mathbb{L}}\neq\emptyset$, train model ${\mathbb{M}}$ on ${\mathbb{L}}$, and compute $\boldsymbol{\delta}$-expansion on the elements of ${\mathbb{L}}$. Otherwise, $\Delta=[\delta_{i}=\delta_{0}]_{i=1}^{b}$.
Repeat /* iterative active learning steps */
  1. (i)
    Active query selection: determine which examples to select for annotation.
  2. (ii)
    Model training: obtain labels and train a deep learning model ${\mathbb{M}}$.
  3. (iii)
    $\boldsymbol{\delta}$-expansion: use ${\mathbb{M}}$ to update $\delta_{i}$ for the recently selected examples.
Until the accumulated budget is exhausted.
We next provide a detailed description of each step.
@@
Initialization: We begin by seeking a suitable embedding space using ${\mathbb{U}}\cup{\mathbb{L}}$, where distances are intended to be inversely correlated with semantic similarity. Self-supervised or representation learning may be used to this end. We choose an initial radius $\delta_{0}$ as described in [39].
(i) Active query selection: In this phase (see Alg 1 below for pseudo-code), the algorithm selects $q$ examples from ${\mathbb{U}}$ to send for labeling, where $q$ denotes the query set size. To this end, it first computes several parameters:
Algorithm 1 DCoM, Active sampling
1:  Input: unlabeled pool ${\mathbb{U}}$, labeled pool ${\mathbb{L}}$, query size $q$, list of ball sizes $\Delta$, trained model ${\mathbb{M}}$ if ${\mathbb{L}}\neq\emptyset$
2:  Output: a set of points to query, and the coverage of ${\mathbb{L}}$
3:  $\forall x\in{\mathbb{U}}$, $M(x)\leftarrow$ normalized margin between the two highest softmax outputs of ${\mathbb{M}}$ if ${\mathbb{L}}\neq\emptyset$, 1 otherwise
4:  Compute $P(C({\mathbb{L}},\Delta))$
5:  Compute ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$
6:  $\delta_{\text{avg}}\leftarrow$ Average($\Delta$)
7:  $G=(V={\mathbb{U}}\cup{\mathbb{L}},E=\{(x,x^{\prime}):x^{\prime}\in B_{\delta_{\text{avg}}}(x)\})$
8:  for $(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta$ do
9:     $\forall x\in B_{\delta_{i}}(x_{i})$: $\forall e=(v,x)\in E$, remove $e$eqparboxCOMMENT# DON’T COVER POINTS TWICE
10:     $\forall e=(x_{i},v)\in E$, remove $e$
11:  end for
12:  ${{\textit{Q}}}\leftarrow\emptyset$
13:  for $i=1,...,q$ do
14:     $\forall x\in{\mathbb{U}}$, $\text{ODR}(x)\leftarrow$Out-Degree Rank
15:     $\forall x\in{\mathbb{U}}$, $R(x)\leftarrow{S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))$ + (1 - ${S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$
16:     $x_{max}\leftarrow\{\text{argmax}_{x\in{\mathbb{U}}}R(x)\}$
17:     ${\textit{Q}}\leftarrow{{\textit{Q}}}\cup\{x_{max}\}$
18:     $\forall x\in B_{\delta_{avg}}(x_{max})$: $\forall e=(v,x)\in E$, remove $e$
19:     $M(x_{max})\leftarrow\text{1}$
20:  end for
21:  return  Q, $P(C({\mathbb{L}},\Delta))$
  • $M(x),~{}\forall x\in{\mathbb{U}}$: the normalized margin between the two highest softmax outputs using the last trained model if ${\mathbb{L}}\neq\emptyset$, 1 otherwise.
  • $\Delta$: $\Delta=[\delta_{i}=\delta_{0}]_{i=1}^{b}$ if ${\mathbb{L}}=\emptyset$, otherwise use $\Delta$ from the last AL iteration.
  • $\delta_{\text{avg}}$: the average of $\Delta$.
  • $P(C({\mathbb{L}},\Delta))$: the probability of coverage of the current ${\mathbb{L}}$ with its corresponding $\Delta$.
  • ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$: the competence score of objective function (6) if ${\mathbb{L}}\neq\emptyset$, 0 otherwise.
  • $\{G_{\delta_{\text{avg}}}=(V,E)\}$: a directed adjacency graph where nodes represent samples and edges connect pairs of nodes if the distance between them in the embedding space is smaller than $\delta_{\text{avg}}$. In order not to cover points more than once, $\forall x_{i}\in{\mathbb{L}}$, $\forall x\in B_{\delta_{i}}(x_{i})$ and $\forall{v}\in{V}$ – prune incoming edges $e=(v,x)\in E$. Additionally, we prune all outgoing edges $e=(x_{i},v)\in E$.
Next, $q$ points are to be selected from ${\mathbb{U}}$ in a greedy manner as follows:
@@
    $\forall x\in{\mathbb{U}}$, compute a ranking score $R(x)={S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))+(1-{S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$.
@@
    Select the vertex $x_{\text{max}}$ with the highest score, $x_{\text{max}}=\text{argmax}_{x\in{\mathbb{U}}}R(x)$.
@@
(ii) Model training:
  1. 1.
    Obtain labels for query set Q, returned in step (i) of Active sampling.
  2. 2.
    Remove set Q from ${\mathbb{U}}$ and add it to ${\mathbb{L}}$.
  3. 3.
    Train the given model ${\mathbb{M}}$ using ${\mathbb{L}}$ (and ${\mathbb{U}}$ if the learner is semi-supervised).
(iii) $\boldsymbol{\delta}$-expansion for new samples. In this step (see Alg 2 below for pseudo-code), we compute a suitable $\delta_{i}$ for each new labeled example $x_{i}\in{\textit{Q}}$, and add it to $\Delta$. First, we use the updated model $M$ to predict labels for all the points in the unlabeled set ${\mathbb{U}}$. We also compute a purity threshold $\tau=m\cdot P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))+b$, where $m$ and $b$ are hyperparameters and $P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))$ the probability of coverage using the old labeled set - has been computed in the first Active sampling step. For each new labeled example $v\in{\textit{Q}}$, we search for the largest radius $\delta_{\text{opt}}$ such that its purity $P\left(\left\{x\in B_{\delta_{\text{opt}}}\left(v\right):f(x)=f(v)\right\}\right)$ is still larger than $\tau$. Assuming that $\pi(B_{\delta}(x))$ a monotonic with $\delta$, we can leverage binary search.
Algorithm 2 DCoM, $\boldsymbol{\delta}$-expansion
1:  Input: unlabeled pool ${\mathbb{U}}$, labeled pool ${\mathbb{L}}$, query pool Q, maximal ball size $\delta_{\text{max}}$, list of ball sizes $\Delta$ for ${\mathbb{L}}\setminus{{\textit{Q}}}$, trained model ${\mathbb{M}}$ from the current iteration and $P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))$ from Active sampling
2:  Output: updated list of ball sizes $\Delta$
3:  $\tau\leftarrow m\cdot P(C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta))+b$
4:  $\hat{{\mathbb{Y}}}\leftarrow$ The predicted label for each $x\in{\mathbb{U}}$ using the model ${\mathbb{M}}$
5:  for $v\in{{\textit{Q}}}$ do
6:     $\delta_{\text{opt}}\leftarrow\text{argmax}_{\delta}[P\left(\left\{x\in B_{\delta}\left(v\right):f(x)=f(v)\right\}\right)>\tau$]
7:     $\Delta\leftarrow\Delta+[\delta_{\text{opt}}]$
8:  end for
9:  return  $\Delta$
Using a sparse representation of the adjacency graph enables DCoM to handle large datasets efficiently without exhaustive space requirements. The algorithm’s complexity, including adjacency graph construction and sample selection, is discussed in App. C.
@@ 4.1 Methodology
This section investigates two deep AL frameworks: a fully supervised approach, wherein a deep model is trained exclusively on the annotated dataset ${\mathbb{L}}$ as a standard supervised task, and a semi-supervised framework trained on both the annotated and unlabeled datasets, ${\mathbb{L}}$ and ${\mathbb{U}}$. Our experimental setup is based on the codebase of [24], ensuring a fair comparison among the various AL strategies by using the same network architectures and sharing all relevant experimental conditions.
More specifically, in the following experiments, we trained ResNet-18 [16] on STL-10 [6], SVHN [25], CIFAR-100 [22] and subsets of ImageNet [9] including ImageNet-50, ImageNet-100, and ImageNet-200 as they appear in [33]. The hyper-parameters we used are detailed in App. A. In the semi-supervised framework, we used FlexMatch [42] and the code provided by [35]. For the experiments, we followed the parameters specified by [42]. The exact parameters are detailed in App. A. While the ResNet-18 architecture may no longer achieve state-of-the-art results on the datasets examined, it serves as a suitable platform to evaluate the efficacy of AL strategies in a competitive and fair environment, where they have demonstrated benefits.
In our comparisons, we used several active learning strategies (or optimization criteria) for selecting a query set: (1) Randomsampling, (2) Min margin— lowest margin between the two highest softmax outputs, (3) Max entropy— highest entropy of softmax outputs [30], (4) Uncertainty— 1 minus maximum softmax output, (5) DBAL[11], (6) CoreSet[28], (7) BALD[21], (8) BADGE[1], (9) ProbCover[39], and (10) LDM-s[4]. When available, we used the code provided in [24] for each strategy. When it was unavailable, we used the code from the repositories of the corresponding papers. Since [4] doesn’t provide code, we compared our results with the corresponding available results reported in their paper, despite differences in the running setup (see App. D.6).
DCoM  and ProbCover require a suitable data embedding. For STL-10 , SVHN, CIFAR-10 and CIFAR-100 we employed the SimCLR [3] embedding. For the ImageNet subsets we used the DINOv2 [26] embedding. Section 4.3 provides an extensive ablation study, which includes alternative embedding spaces. In all experiments, identical settings and hyper-parameters are used for our method, as detailed in App. A. The robustness of our method to these choices is discussed in App. B.
@@ 4.2 Main results
Figure 1: Mean accuracy difference between several AL algorithms and random queries. Positive mean difference indicates that the corresponding AL method is beneficial. The errors bars correspond to standard error, based on 5-10 repetitions (3 for the ImageNet subsets). While some methods perform well only in specific budget regimes, DCoM consistently achieves the best results across all budgets.
Fully supervised framework. We evaluate various AL methods by training a ResNet-18 model from scratch for each different AL strategy, using the labels of the queries selected by the respective method. In each AL iteration, a new model is trained from scratch using the updated labeled set ${\mathbb{L}}$. This process is repeated for several AL iterations. Fig. 1 shows the difference between the accuracy obtained by each AL method, and the accuracy obtained when training a similar network with a random query set of the same size (the baseline of no active learning), for each AL iteration. The final accuracy of each method is shown in Tables 1-2 for the ImageNet-100 and CIFAR-100 datasets, as mean $\pm$ Standard Error (STE). Additional datasets are shown in App. D.
Table 1: Model final accuracy when varying the size of the labeled set ${\mathbb{L}}$ (row) and the respective AL strategy (column), using the ImageNet-100 dataset
Table 2: Same as Table 1, using the CIFAR-100 dataset.
@@
Figure 2: Comparison of AL strategies in a semi-supervised learning framework. Each bar represents the mean test accuracy over 3 repetitions of FlexMatch, trained using a total of 400 (left) and 1,000 (right) labeled examples on the CIFAR-100 dataset (an average of 4 and 10 labeled examples per class respectively). Three rounds of AL are considered, selecting 100 examples in the first round, 300 in the second, and 600 in the third.

### v3

@@ 3 Our method: DCoM
[…]
(a) Initial active sampling using $\delta_{0}$ in DCoM, following ProbCover approach.
(b) DCoM’s updated $\Delta$ list, which includes a suitable $\delta$ value for each example in $\mathbb{L}$}
Figure 2: Illustration of Step 4: $\Delta$-Adjustment in DCoM, comparing the coverage obtained by DCoM (b) with ProbCover (a) using budget $b=15$. Selected points are marked by x, and the covered area is shaded in light gray. While ProbCover maintains a constant $\delta$ for each example, DCoM seeks to maximize coverage while preserving the purity of the balls by fitting a suitable radius for each center.
3.1 Framework and definitions
$\forall x\in{\mathbb{X}}$, let $M(x)$ denote the normalized margin between the two highest softmax outputs from the last trained model $\mathbb{M}$}. Define $\Delta$ as the set of individual $\delta$-values for each labeled example in $\mathbb{L}$}. As in the theoretical analysis, $P[C({\mathbb{L}},\Delta)]$ denotes the probability of coverage for $\mathbb{L}$} given $\Delta$. The competence score ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$ is defined by the objective function (5). The graph $\{G_{\delta}=(V,E)\}$ is a directed adjacency graph where nodes are samples, and edges connect pairs of nodes if their distance in the embedding space is less than $\delta$.
@@ 3.2 Active learning method
Active Learning (AL) is performed iteratively (see Fig. 1): at each iteration $i$, a query set Q of $q$ unlabeled examples is selected based on a specific strategy, where $q$ denotes the query set size and Q denotes the current query set. These examples are then labeled by an oracle, added to the labeled set $\mathbb{L}$}, and removed from the unlabeled pool $\mathbb{U}$}. Subsequently, the algorithm parameters may be updated as needed, depending on the specific requirements of the algorithm. This process is repeated until the label budget is exhausted or predefined termination conditions are met.
Our AL method DCoM involves several distinct steps:
@@
Step 1: Initialization. Identify an appropriate embedding space for ${\mathbb{U}}\cup{\mathbb{L}}$, where distances are expected to be inversely correlated to semantic similarity. This can be achieved using self-supervised or representation learning techniques. Select an initial radius $\delta_{0}$ as outlined in (Yehuda et al. 2022), and initialize the parameter $\Delta=\emptyset$. If $\mathbb{L}$$\neq\emptyset$, skip Step 2 for now and begin the iterations with Step 3.
Step 2: Active sampling. Select $q$ examples from $\mathbb{U}$} for labeling (see Alg 1 for the pseudo-code). Begin by computing the normalized margin $M(x)$ $\forall x\in{\mathbb{U}}$ using the latest model $\mathbb{M}$}. Next, construct a directed adjacency graph $\{G_{\delta_{\text{avg}}}=(V,E)\}$, where $\delta_{\text{avg}}$ is the average of $\Delta$, or $\delta_{0}$ if $\Delta$ is empty. To avoid covering points multiple times, $\forall x_{i}\in{\mathbb{L}}$, $\forall x\in B_{\delta_{i}}(x_{i})$ and $\forall v\in{V}$, prune all incoming edges $e=(v,x)\in E$. Additionally, prune all outgoing edges $e=(x_{i},v)\in E$.
Using this graph, compute the Out-Degree-Rank (ODR) $\forall x\in{\mathbb{U}}$. Likewise, compute the probability of coverage $P[C({\mathbb{L}},\Delta)]$ of the current $\mathbb{L}$} and its corresponding $\Delta$, and use these quantities to determine ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$. Now select $q$ points from $\mathbb{U}$} in a greedy manner as follows:
@@
    $\forall x\in{\mathbb{U}}$, compute a ranking score
    $\hskip-17.07182ptR(x)={S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))+(1-{S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x).$
@@
    Select the vertex $x_{\text{max}}$ with the highest score
    $x_{\text{max}}=\text{argmax}_{x\in{\mathbb{U}}}R(x).$
@@
Algorithm 1 DCoM, Active sampling
Input: unlabeled pool $\mathbb{U}$}, labeled pool $\mathbb{L}$}, query size $q$, list of ball sizes $\Delta$, trained model $\mathbb{M}$} if $\mathbb{L}$$\neq\emptyset$
Output: a set of points to query, and the coverage of $\mathbb{L}$}
[…]
2: Compute $P[C({\mathbb{L}},\Delta)]$
3: Compute ${S_{\mathcal{L}}}({\mathbb{L}},\Delta)$
[…]
6:for $(x_{i},\delta_{i})\in{\mathbb{L}}\times\Delta$ do
7:  $\forall x\in B_{\delta_{i}}(x_{i})$: $\forall e=(v,x)\in E$, remove $e$
8:  $\forall e=(x_{i},v)\in E$, remove $e$
9:end for
[…]
11:for i=1,,qi=1,...,q do
12:  $\forall x\in{\mathbb{U}}$, $\text{ODR}(x)\leftarrowOut$-Degree Rank
13:  $\forall x\in{\mathbb{U}}$, $R(x)\leftarrow{S_{\mathcal{L}}}({\mathbb{L}},\Delta)\cdot(1-M(x))$ + (1 - ${S_{\mathcal{L}}}({\mathbb{L}},\Delta))\cdot\text{ODR}(x)$
14:  $x_{max}\leftarrow\{\text{argmax}_{x\in{\mathbb{U}}}R(x)\}$
15:  ${\textit{Q}}\leftarrow{{\textit{Q}}}\cup\{x_{max}\}$
16:  $\forall x\in B_{\delta_{avg}}(x_{max})$: $\forall e=(v,x)\in E$, remove $e$
17:  $M(x_{max})\leftarrow\text{1}$
18:end for
19:return Q, $P[C({\mathbb{L}},\Delta)]$
Step 3: Model training: Obtain labels for the query set Q derived in Step 2. Remove Q from $\mathbb{U}$} and add it to $\mathbb{L}$}. Finally, train model $\mathbb{M}$} using the updated $\mathbb{L}$} (and also $\mathbb{U}$} if the learner employs semi-supervised learning).
Step 4: $\boldsymbol{\Delta}$-Adjustment. Determine a suitable $\delta_{i}$ for each new labeled example $x_{i}\in{\textit{Q}}$ and add it to $\Delta$, see Alg 2 for pseudo-code and illustration in Fig. 2. Begin by using the updated model $\mathbb{M}$} to predict labels for all points in the unlabeled set $\mathbb{U}$}. Likewise, compute a purity threshold $\tau=m\cdot P[C({\mathbb{L}}\setminus{\textit{Q}},\Delta)]+d$, where [^2] The linear transformation defined by $m,d$ adjusts the range of coverage scores into the relevant range of purity scores $[b,b+m]$. $P[C({\mathbb{L}}\setminus{\textit{Q}},\Delta)]$, the probability of coverage using the old labeled set, is derived in Step 2. Finally, for each new labeled example $v\in{\textit{Q}}$, search for the largest radius $\delta_{\text{opt}}$ whose purity $P\left(\left\{x\in B_{\delta_{\text{opt}}}\left(v\right):f(x)=f(v)\right\}\right)$ remains above $\tau$. Assuming that $\pi(B_{\delta}(x))$ is monotonic with $\delta$, use binary search.
Algorithm 2 DCoM, $\boldsymbol{\Delta}$-Adjustment
Input: unlabeled pool $\mathbb{U}$}, labeled pool $\mathbb{L}$}, query pool Q, maximal ball size $\delta_{\text{max}}$, list of ball sizes $\Delta$ for ${\mathbb{L}}\setminus{{\textit{Q}}}$, trained model $\mathbb{M}$} from the current iteration and $P[C({\mathbb{L}}\setminus{{\textit{Q}}},\Delta)]$ from Step 2
Output: updated list of ball sizes $\Delta$
[…]
3:for $v\in{{\textit{Q}}}$ do
4:  $\delta_{\text{opt}}\leftarrow\text{argmax}_{\delta}[P\left(\left\{x\in B_{\delta}\left(v\right):f(x)=f(v)\right\}\right)>\tau$]
5:  $\Delta\leftarrow\Delta+[\delta_{\text{opt}}]$
6:end for
7:return $\Delta$
@@ 4.1 Methodology
This section investigates two deep AL frameworks: a fully supervised approach, wherein a deep model is trained exclusively on the annotated dataset $\mathbb{L}$} as a standard supervised task, and a semi-supervised framework trained on both the annotated and unlabeled datasets, $\mathbb{L}$} and $\mathbb{U}$}. Our experimental setup is based on the codebase of (Munjal et al. 2022), ensuring a fair comparison among the various AL strategies by using the same network architectures while sharing all the relevant experimental conditions.
More specifically, we trained ResNet-18 (He et al. 2016) on the following benchmark datasets: STL-10 (Coates, Ng, and Lee 2011), SVHN (Netzer et al. 2011), CIFAR-100 (Krizhevsky, Hinton et al. 2009) and subsets of ImageNet (Deng et al. 2009) including ImageNet-50, ImageNet-100, and ImageNet-200 as defined in (Van Gansbeke et al. 2020). The hyper-parameters are detailed in Appendix A. In the semi-supervised framework, we used FlexMatch (Zhang et al. 2021) and the code provided by (Wang et al. 2022), adopting the parameters specified by (Zhang et al. 2021) (see details in Appendix A). While the ResNet-18 architecture may no longer achieve state-of-the-art results on the datasets examined, it serves as a suitable platform to evaluate the efficacy of AL strategies in a competitive and fair environment, where they have demonstrated benefits.
In our comparisons, we used several active learning strategies for selecting a query set: (i) Randomsampling, (ii) Min margin— lowest margin between the two highest softmax outputs; (iii) Max entropy— highest entropy of softmax outputs; (iv) Uncertainty— 1 minus the highest softmax output; (v) DBAL(Gal, Islam, and Ghahramani 2017); (vi) CoreSet(Sener and Savarese 2018); (vii) BALD(Kirsch, Van Amersfoort, and Gal 2019); (viii) BADGE(Ash et al. 2019); (ix) ProbCover(Yehuda et al. 2022); and (x) LDM-s(Cho et al. 2023). When available, we used the code provided in (Munjal et al. 2022) for each strategy. When it was unavailable, we used the code from the repositories of the corresponding papers. Since (Cho et al. 2023) doesn’t provide code, we compared our results with the corresponding available results reported in their paper, despite differences in the running setup (see Appendix E.5).
DCoM and ProbCover require a suitable data embedding. For STL-10 , SVHN, CIFAR-10 and CIFAR-100 we employed the SimCLR (Chen et al. 2020) embedding. For the ImageNet subsets we used the DINOv2 (Oquab et al. 2023) embedding. Section 4.3 provides an extensive ablation study, which includes alternative embedding spaces. In all experiments, identical settings and hyper-parameters are used as detailed in Appendix A. The robustness of our method to these choices is discussed in Appendix C.
Using a sparse representation of the adjacency graph enables DCoM to handle large datasets efficiently without exhaustive space requirements. The algorithm’s complexity, including adjacency graph construction and sample selection, is discussed in Appendix D.
@@ 4.2 Main results
Figure 3: Mean accuracy difference between AL algorithms and random queries. Positive mean difference indicates that the corresponding AL method is beneficial. The errors bars correspond to standard error, using 5-10 repetitions (3 for the ImageNet subsets). While some methods perform well only in specific budget regime, DCoM consistently achieves the best results across all budgets. Cumulative budget is evenly spaced for easy range comparison.
Fully supervised framework. We evaluate various AL methods by training a ResNet-18 model from scratch for each different AL strategy, using the labels of the queries selected by the respective method. In each AL iteration, a new model is trained from scratch using the updated labeled set $\mathbb{L}$}. This process is repeated for several AL iterations. Fig. 3 shows the difference between the accuracy obtained by each AL method, and the accuracy obtained when training a similar network with a random query set of the same size (the baseline of no active learning), for each AL iteration. The final accuracy of each method is shown in Table 1 for the CIFAR-100 dataset, as mean $\pm$ Standard Error (STE). Similar tables for the remaining datasets are shown in Appendix E.
@@
Table 1: Model accuracy with varying labeled set sizes $\mathbb{L}$} (row) and AL strategies (column) while training directly on CIFAR-100 images.
@@
Figure 4: Comparison of AL strategies in a semi-supervised learning framework. Each bar shows the mean test accuracy over 3 repetitions of FlexMatch, trained with 400 and 1,000 labeled examples on CIFAR-100. Three AL rounds are conducted, selecting 100, 300, and 600 examples in each round.